# Comparative analysis of diverse methodologies for portfolio optimization leveraging quantum annealing techniques

### 05 Mar 2024 v1

@@ II Portfolio optimization formulation
Portfolio optimization is a challenging problem in finance and management. Its objective is to find a balance between risk and return. The formulation of portfolio optimization depends on the given market conditions. In this work, we study the most general formulation given by Markowitz’s mean-variance model:
$\max_{x}\left\{\sum_{i}^{n}\mu_{i}x_{i}-q\sum_{i,j}^{n}\sigma_{ij}x_{i}x_{j}-\lambda\left(B-\sum_{i}^{n}x_{i}\right)\right\},$ (1)
where $\mu_{i}$ is the expected return of assets, $\sigma_{ij}$ is the risk of assets, and $B$ is the full budget. To solve the portfolio optimization problem with quantum annealing, it is necessary to transform the formulation into an Ising model through various conversions, such as binary, unary, or sequential encoding. [89, 65]. The given formulation in Markowitz’s mean-variance model is a binary constrained mean-variance model. This model addresses the binary portfolio optimization problem, where the variables are limited to values of 0 or 1. To adapt the portfolio optimization problem for quantum annealing, we convert it into a Quadratic Unconstrained Binary Optimization (QUBO) problem. The problem is encoded into a problem Hamiltonian $H_{P}$, which is defined by pauli matrices
@@
Quantum-based algorithms have demonstrated their efficacy in enhancing portfolio optimization outcomes, such as the works using gate-model Quantum computing in [59, 64, 60, 91, 62, 61, 92, 93] and the ones using Quantum annealing algorithms in [65, 67, 68, 69, 70, 70, 94, 27, 73, 74, 75, 95]. Despite the limited connectivity of quantum annealers, they provide more accessible stable qubits than gate-model quantum computer in the NISQ era. Consequently, our focus in this study will be on evaluating the performance of different quantum annealing algorithms. To make a comprehensive comparison of QA algorithms for solving the portfolio optimization problem, we transform the portfolio optimization model into an ising model and simulate the annealing process under the guidance of time dependent Schrodinger equation. Various adaptations of QA algorithms are studied and compared to provide guidance for their implementation on real quantum hardware. Furthermore, the reverse annealing of each algorithm is investigated and introduced in the following sections.
@@ III.4 CD-QA
In standard QA, the ground state of $H_{P}$ can be found with high probability after annealing time that relates to the minimum energy gap. In some problems with exponentially small energy gap, the required annealing time can be extremely long. To make quantum annealing more practical to use, the idea of suppressing nonadiabatic transition during short time evolution was proposed [84, 85, 86], the non-adiabatic technique is named as counter-diabatic quantum annealing (CD-QA). The effectiveness of CD-QA has been proven in the works [100, 101, 102, 103, 104, 105, 106, 107, 108, 109, 110, 111, 27, 112, 113, 114, 115]

### 11 Mar 2024 v2

@@ II Portfolio optimization formulation
Portfolio optimization is a challenging problem in finance and management. Its objective is to find a balance between risk and return. For instance, the general formulation of Markowitz’s mean-variance model can be expressed as follows [59]:
$\max_{x}\left\{\sum_{i}^{n}\mu_{i}x_{i}-q\sum_{i,j}^{n}\sigma_{ij}x_{i}x_{j}-\lambda\left(B-\sum_{i}^{n}x_{i}\right)^{2}\right\},$ (1)
the given formulation in Markowitz’s mean-variance model is a binary constrained mean-variance model. This model addresses the binary portfolio optimization problem, where the variables are limited to values of 0 or 1, $\mu_{i}$ is the expected return of assets, $\sigma_{ij}$ is the risk of assets, and $B$ is the full budget.
The formulation of a particular portfolio optimization depends on the given market conditions. To adapt a portfolio optimization problem for quantum annealing, it is necessary to transform the formulation into a Quadratic Unconstrained Binary Optimization (QUBO) problem through various conversions, such as binary, unary, or sequential encoding. [89, 65]. The problem is encoded into a problem Hamiltonian $H_{P}$, which is defined by pauli matrices
@@
Quantum-based algorithms have demonstrated their efficacy in enhancing portfolio optimization outcomes, such as the works using gate-model Quantum computing in [59, 64, 60, 91, 92, 62, 61, 93, 94] and the ones using Quantum annealing algorithms in [65, 67, 68, 69, 70, 70, 95, 27, 73, 74, 75, 96]. Despite the limited connectivity of quantum annealers, they provide more accessible stable qubits than gate-model quantum computer in the NISQ era. Consequently, our focus in this study will be on evaluating the performance of different quantum annealing algorithms. To make a comprehensive comparison of QA algorithms for solving the portfolio optimization problem, we transform the portfolio optimization model into an ising model and simulate the annealing process under the guidance of time dependent Schrodinger equation. Various adaptations of QA algorithms are studied and compared to provide guidance for their implementation on real quantum hardware. Furthermore, the reverse annealing of each algorithm is investigated and introduced in the following sections.
@@ III.4 CD-QA
In standard QA, the ground state of $H_{P}$ can be found with high probability after annealing time that relates to the minimum energy gap. In some problems with exponentially small energy gap, the required annealing time can be extremely long. To make quantum annealing more practical to use, the idea of suppressing nonadiabatic transition during short time evolution was proposed [84, 85, 86], the non-adiabatic technique is named as counter-diabatic quantum annealing (CD-QA). The effectiveness of CD-QA has been proven in the works [101, 102, 103, 104, 105, 106, 107, 108, 109, 110, 111, 112, 27, 113, 114, 115, 116, 117]